OpenStudy (anonymous):

i need help with verifying cos(5x)/sin(x)+cos(5x)/cos(x)=cos(4x)/sin(x)cos(x)

OpenStudy (loser66):

re check the problem, I think you wrote something not correct. ( from the numerator)

OpenStudy (anonymous):

I double checked .. its the same

OpenStudy (anonymous):

what about this one tan^-1(x)+tan^-1(1/x)=pi/2, x>0

OpenStudy (loser66):

I don't know, both :( I am sorry.

OpenStudy (anonymous):

thanks for the thought

OpenStudy (anonymous):

and this one cos^-1(x)+cos^-1(-x)=pi

OpenStudy (jdoe0001):

can you post a quick screenshot of the material?

OpenStudy (jdoe0001):

do you know how to take a screenshot?

OpenStudy (jdoe0001):

well.. anyhow ... \(\bf \cfrac{cos(5x)}{sin(x)}+\cfrac{cos(5x)}{cos(x)}=\cfrac{cos(4x)}{sin(x)cos(x)}\\\implies \cfrac{2cos(5x)}{sin(x)cos(x)}=\cfrac{cos(4x)}{sin(x)cos(x)} \) withouth going further... I did a quick check if they match up.... so check the graph below here

OpenStudy (jdoe0001):

anyhow.... in the end, they won't match, thus the identities aren't equal

OpenStudy (anonymous):

thank you very much , did u check the others ?

OpenStudy (anonymous):

this cos^-1(x)+cos^-1(-x)=pi and this tan^-1(x)+tan^-1(1/x)=pi/2, x>0 ???

OpenStudy (jdoe0001):

haven't.... need to dash in a few secs :(